### v1

# Generative Bid Shading in Real-Time Bidding Advertising

Abstract.
Bid shading plays a crucial role in Real-Time Bidding (RTB) by adaptively adjusting the bid to avoid advertisers overspending. Existing mainstream two-stage methods, which first model bid landscapes and then optimize surplus using operations research techniques, are constrained by unimodal assumptions that fail to adapt for non-convex surplus curves and are vulnerable to cascading errors in sequential workflows. Additionally, existing discretization models of continuous values ignore the dependence between discrete intervals, reducing the model’s error correction ability, while sample selection bias in bidding scenarios presents further challenges for prediction. To address these issues, this paper introduces Generative Bid Shading (GBS), which comprises two primary components: 1) an end-to-end generative model that utilizes an autoregressive approach to generate shading ratios by stepwise residuals, capturing complex value dependencies without relying on predefined priors; and 2) a reward preference alignment system, which incorporates a channel-aware hierarchical dynamic network (CHNet) as the reward model to extract fine-grained features, along with modules for surplus optimization and exploration utility reward alignment, ultimately optimizing both short-term and long-term surplus using group relative policy optimization (GRPO). Extensive experiments on both offline and online A/B tests validate GBS’s effectiveness. Moreover, GBS has been deployed on the Meituan DSP platform, serving billions of bid requests daily.
Bid Shading, Generative Model, Auto-bidding
isbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Computational advertising
1. Introduction
Online advertising is now a primary way for companies to attract users and build their brands. Real-Time Bidding (RTB) is the key system in programmatic advertising, allowing auctions in milliseconds. When a user visits a website, the Supply-Side Platform (SSP) sends an ad request to the Ad Exchange (ADX), where Demand-Side Platforms (DSPs) bid based on request information. The highest bidder wins the chance to show their ad. The auction way has changed a lot due to transparency issues in auctions and compatibility challenges with emerging technologies (AppNexus, 2017; Getintent, 2017; Sluis, 2017). Most major platforms, following Google (Bigler, 2019), have switched from Second-Price Auctions (SPA) to First-Price Auctions (FPA). In SPA, advertisers usually bid through their actual value, but in FPA or non-ideal SPA settings, this can cause them to pay much more than necessary. For example, if an advertiser bids \$5 while the second-highest bid is only \$2, they would overpay up to \$3. By dynamically adjusting bids toward the minimum winning price, bid shading (Crespi and Sexton, 2005) significantly optimizes cost structures and directly improves resource allocation efficiency in the trillion-dollar digital advertising market.
Figure 1. Overview of two-stage bid shading process.
Bid shading is a censored problem (Huang et al., 2024) because of the inaccessibility of the minimum winning price. Some methods assume an uncensored scenario in which the ADX will return the minimum winning price regardless of the auction outcome, allowing them to fit an optimal shading ratio using linear regression. However, this assumption is unrealistic in real-world DSP environments: in FPA, the minimum winning price is unobservable regardless of whether the bid wins or loses, resulting in no ground truth for bid shading. Most existing bid shading strategies adopt a two-stage framework, as illustrated in Fig. 1. The first stage is machine learning (ML), which estimates the bid landscape (i.e., the winning price or winning rate distribution) to compute the winning rate and expected cost for varying bids. The second stage is operations research (OR), which employs optimization algorithms (e.g., bisection search, golden section search) to identify the bid that maximizes surplus, using the ML predictions as input.
Figure 2. The winning rate and surplus curves derived from Meituan’s non-ideal SPA data show that surplus first increases and then decreases as bids increase, but do not strictly follow a standard unimodal pattern. As a result, two-stage methods are challenging in finding the global optimum. In contrast, GBS takes a broader view of the generative space, allowing it to escape the local optimum effectively.
Despite its widespread adoption, the two-stage method faces two significant limitations. First is the non-convex surplus profile: existing methods typically strongly assume that the surplus optimization curve is unimodal (Pan et al., 2020; Zhou et al., 2021), thereby enabling search algorithms to determine the optimal shading ratio. However, in real-world scenarios, winning rate curves are often non-smooth, and cost curves in non-ideal SPA exhibit significant fluctuations. This results in a non-convex surplus profile, causing optimization algorithms to converge to local optima, as illustrated in Fig. 2 (a). Second, cascading errors from the ML+OR pipeline: the sequential workflow introduces cascading errors. In the precise mathematical operations of the OR stage, errors originating from the ML stage can be further amplified. Consequently, even theoretically optimal OR algorithms may fail to achieve practical optimality. An ideal approach is to evaluate all bid candidates in real time to identify the globally optimal bid. However, such an exhaustive search is computationally infeasible in a millisecond-latency RTB system.
A recent study (Gong et al., 2023) has attempted to generate the shading ratio in an end-to-end manner. However, their use of negative log-surplus as the loss function in supervised learning encounters the same limitations as two-stage methods when faced with non-convex surplus curves: convergence to local optima. Furthermore, their simple discriminative models struggle to meet the high-precision requirements of bidding scenarios. Traditional generative algorithms, such as conditional variational autoencoders (CVAE (Sohn et al., 2015)), are constrained by predefined prior distributions (e.g., Gaussian), which limits their performance ceiling in shading ratio generation. To overcome distributional assumptions and mitigate outlier interference, existing direct scalar-output methods often discretize continuous values (Sun et al., 2024; Lin et al., 2023). However, this approach fails to leverage conditional dependencies between discrete intervals; discretization is only reflected in label definitions without mutual interdependence (e.g., 0.3 is closer to 0.4 than 0.2), and the lack of error-correction mechanisms leads to suboptimal results. Additionally, DSPs optimize holistic revenue during bidding, causing model performance to be heavily influenced by bid price distributions. This introduces significant data selection bias, as the training data fails to represent the real action space of bids.
In this paper, we propose a novel end-to-end Generative Bid Shading (GBS) framework that leverages an autoregressive generative model and reward-preference alignment to address the challenges above from a global perspective, as illustrated in Fig. 2 (b). GBS comprises two main components:
1). An autoregressive generative model: Inspired by the success of large language models (LLMs) in recommendation and regression tasks, we adopt an encoder-decoder autoregressive architecture to generate shading ratios. We construct a data-driven vocabulary of primitive tokens, encoding any shading ratio scalar as a token sequence. The model decomposes shading ratio generation into a conditional and sequential modeling process, where each step predicts a ratio component, and the output at each time step serves as input for the next. This method effectively captures dependencies across multiple prediction steps and progressively approximates the shading ratio in a residual manner. The token sequence generation enables coverage of a broader value range and adaptation to complex distributions. To accelerate model convergence and ensure stability, we employ teacher forcing during pre-training and introduce gumbel softmax to guarantee gradient backpropagation and consistency between training and inference.
2). A reward preference alignment system: To overcome the limitations of traditional approaches, we introduce a preference alignment reward system and train the model using reinforcement learning in the generated space. We first propose a channel-aware hierarchical dynamic network (CHNet) as the reward model to capture fine-grained features for precise bid landscape modeling. Next, we introduce two preference alignment modules to guide learning. Specifically, surplus optimized alignment maximizes the final surplus by ensuring that bids for obtainable requests are not overbidding, while exploration utility alignment incorporates bid exploration into the bid shading system, alleviating sample selection bias and optimizing long-term surplus. Finally, we employ the group relative policy optimization (GRPO) for reinforcement learning to achieve efficient and stable policy updates.
In summary, our contributions are as follows:
  • We identify the suboptimality in two-stage bid shading methods and propose a generative framework that autoregressively generates shading ratios with error correction capabilities from a global perspective. To the best of our knowledge, this work represents the first implementation of generative bid shading in large-scale online systems.
  • We design a CHNet that captures fine-grained features for the shifting bidding landscape. Based on this, we introduce two reward preference alignment modules: one maximizes immediate surplus by avoiding overbidding; the other mitigates data selection bias via strategic bid exploration. This paradigm enhances generalization across diverse business scenarios and extensibility.
  • Through extensive experiments on both offline and online A/B tests, we show that our GBS significantly outperforms the state-of-the-art methods. GBS has been deployed in Meituan DSP and serves billions of bid requests daily, demonstrating industrial viability and efficiency.
2. Related Work
Bid Landscape Forecasting. Bid landscape Forecasting is a crucial component of RTB advertising systems, primarily aiding in forecasting the winning price distribution of bid requests. Early research often assumed that winning price followed specific probability distributions, such as Gaussian (Wu et al., 2015), Log-Normal (Cui et al., 2011), Gamma (Chapelle, 2015; Zhu et al., 2017), or other parametric distributions (Ren et al., 2017, 2016). To overcome the limitations of these predefined distributional assumptions, subsequent studies introduced distribution-free models (Ren et al., 2019; Huang et al., 2024; Li et al., 2022). However, existing methods generally overlook the modeling of fine-grained features. This paper proposes a channel-aware hierarchical dynamic network to achieve fine-grained modeling as a foundation for the reward system.
Bid Shading. Bid shading helps advertisers avoid overpaying. Existing methods fall into three categories: 1) Direct regression of the shading ratio (Gligorijevic et al., 2020), which predicts the optimal ratio but unrealistically assumes access to all winning prices; 2) Two-stage ML + OR optimization (Pan et al., 2020; Zhou et al., 2021; Si et al., 2023), where machine learning estimates bid landscape and operations research maximizes surplus; and 3) Non-parametric estimation (Zhang et al., 2021), which uses dynamic binning without parametric models, but still needs optimal shading labels and cannot fully capture individual differences.
Generative Model and Preference Alignment. With the rapid development of large language models (LLMs), generative models are increasingly used in areas such as generative recommendation (Deng et al., 2025) and search (Guo et al., 2025). Preference learning, including Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), aligns model outputs with human preferences but is inefficient. Direct Preference Optimization (DPO) (Rafailov et al., 2023) and its variants (Azar et al., 2024; Chowdhury et al., [n. d.]) address this by directly adjusting model policies using preference data. However, they face challenges like sample sparsity in recommendation and advertising. Group Relative Preference Optimization (GRPO) (Shao et al., 2024) uses explicit reward functions and group-based evaluation to enhance decision-making and training efficiency. In this paper, we present the first exploration of applying the generative model to bid shading, leveraging reward systems to guide the model beyond the constraints of traditional methods.
3. Preliminaries
3.1. Problem Definition
Suppose each DSP has an online advertising system to process bidding requests in the RTB scenario. When users visit web pages containing advertising opportunities, each DSP selects the candidate advertisement with the highest advertising value and places a bid. Once the auction ends, the winning advertisement is displayed, and the win/loss tags are returned to the DSP. Let $X=\left\{x_{1},x_{2},\ldots,x_{N}\right\}$ represent the bid requests, where $x_{i}$ includes features such as user and contextual information. The value of each request is calculated through metrics such as the click-through rate (CTR) and conversion rate (CVR), denoted as $v=\left\{v_{1},v_{2},\ldots,v_{N}\right\}$, which signifies the expected value derived from an advertisement request and serves as the pre-shadow bid [^1] This step usually solves an optimization problem with constraints and will not be elaborated on in detail in this paper. In the bid shading system, we determine a shading ratio $\alpha_{i}$ for each request using a generative model, and multiply it with the unshaded bid to obtain the final bid $B=\left\{v_{1}*\alpha_{1},v_{2}*\alpha_{2},\ldots,v_{N}*\alpha_{N}\right\}$. The expected surplus of a bid request $x_{i}$ at the bid price $b_{i}$ is $surplus(x_{i},b_{i})=(v_{i}-c_{i})\cdot Wr(b_{i}\mid x_{i})$, where $Wr(\cdot)$ represents the winning rate and $c_{i}$ represents the cost incurred if this request is won, $c_{i}=b_{i}$ in FPA scenario. For the bid shading system, the bid prices $b^{*}_{i}$ maximize the expected surplus:
(1) $b^{*}_{i}=\underset{b_{i}\in(b_{min},b_{max})}{\arg\max}(v_{i}-c_{i})\cdot Wr(b_{i}\mid x_{i}),$
where $b_{min}$ and $b_{max}$ are manually set as the lower and upper limits for bid shading. The DSP initiates a bid using the final $b^{*}_{i}$ and compares it against the bids from other advertisers. Once the auction concludes, the win/lose label $y^{*}_{i}$ and the final cost $c_{i}$ are determined. Suppose $z_{i}$ represents the winning price, and we have
(2) $y_{i}^{*}=\begin{cases}1,&b_{i}^{*}>z_{i},\\ 0,&b_{i}^{*}\leq z_{i}.\end{cases}$
It is important to note that $z_{i}$ is not disclosed to the DSP. The actual surplus of the bid request $x_{i}$ at $b_{i}^{*}$ is ${surplus}(x_{i},b_{i}^{*})=(v_{i}-c_{i})\cdot y_{i}^{*},$ which serves as a crucial evaluation metric for bid shading.
3.2. Generative Model Preprocessing
Many methods that output a scalar by employing discretization of continuous results and convert them into classification tasks to simplify the learning process (Sun et al., 2024; Lin et al., 2023). However, predictions for the discretized buckets are typically generated independently, limiting opportunities for effective error correction. Current generative recommendation methods frequently utilize residual quantization (RQ-VAE) and hierarchical k-means clustering to obtain index IDs, and generate ID sequences through an autoregressive model (Rajput et al., 2023). In addition, generative regression (Ma et al., 2024) also uses this way to model scalar generation, achieving excellent performance. Inspired by these methods, we propose generating the shading ratio using an autoregressive encoder-decoder architecture.
Specifically, we introduce a vocabulary $\mathcal{V}=\left\{w_{j}\right\}_{j=1}^{V}$ where each element $w_{i}$ represents a predefined scalar value, analogous to tokens in the language model. The vocabulary embedding matrix is denoted as ${E}\in\mathbb{R}^{V\times D}$, where $V$ consists of the vocabulary size, and $D$ is the dimension of the token embeddings. Each shading ratio can be encoded as a sequence of tokens $s_{i}=\left\{s_{i}^{1},\ldots,s_{i}^{L}\right\}$, with $L$ indicating the length of the sequence. This process is known as tokenization. In contrast to generative recommendation systems that index directly based on ID, we propose a label decoding function $g(\cdot)$ to determine the actual scalar value from the token ID output by the model, thereby deriving the shading factor $\alpha_{i}=g(s_{i})=\sum_{l=1}^{L}g(s_{i}^{l})\in\mathbb{R}$.
Figure 3. The proposed generative model employs an encoder-decoder architecture to predict the token sequence autoregressively. It determines the specific value of each token using a decoding function $g(\cdot)$, then sums these values to obtain the final shading ratio. Additionally, it utilizes the gumbel softmax and teacher forcing strategies to ensure stability and accelerate convergence.
4. Method
In this section, we provide a detailed explanation of the proposed GBS. First, we introduce the end-to-end autoregressive generative model, which maps the shading factor through the output token sequence. Next, we present the reward preference alignment system, which employs a channel-aware hierarchical dynamic network as the reward model and utilizes GRPO to guide policy learning via two reward alignment modules.
4.1. End-to-End Generative Model
Our generative model decomposes the final task into multiple classification subtasks, utilizing inter-step dependencies to approximate the final shading ratio accurately. This framework allows each prediction step the flexibility to choose from the vocabulary, and the expanded solution space enables the model to generate a broader spectrum of potential sequences.
Tokenizer. The vocabulary $\mathcal{V}$ comprises predefined ratio slots as tokens, allowing the model to generate shading ratios that closely approximate actual values. Two key properties are prioritized when constructing $\mathcal{V}$: 1) Each token must be unique, enabling $\mathcal{V}$ to represent all possible shading ratios with a finite token set; 2) Token frequencies should be relatively uniform to prevent class imbalance. We employ the dynamic percentile-based algorithm (Ma et al., 2024) for automated vocabulary construction from specified datasets. It accelerates tail-value reduction, rapidly decreases inter-update variance, and enhances algorithmic reliability (see Appendix A for details).
Due to the lack of ground-truth labels for shading ratios, we construct the vocabulary $\mathcal{V}$ using the shading ratio from the two-stage method and perform tokenization, which involves transforming ratios $\left\{\alpha_{i}\right\}_{i=1}^{N}$ into discrete tokens $\{s_{i}=\{s_{i}^{1},\ldots,s_{i}^{L}\}\}_{i=1}^{N}$. To reduce the complexity of learning encoded labels, the tokenized sequences should adhere to three principles: 1) The reconstructed value $g(s_{i})$should minimize error relative to the real ratio; 2) The sequence length $L$ of $s_{i}$ should be minimized; 3) The sequence should maintain monotonicity: $s_{i}^{1}\geq s_{i}^{2}\geq\ldots\geq s_{i}^{L}$. Sequences that meet these conditions enable the model to incrementally output residual results, thereby significantly simplifying the learning process. Consequently, we employ a greedy strategy during tokenization, generating training data for the pre-training stage.
Encoder-Decoder Architecture. As illustrated in Fig. 3, our model employs a Transformer-based architecture, consisting of an embedding layer, an encoder for modeling features, and a decoder for generating shading ratios. Specifically, the embedding layer maps features through the parameter matrix. For dense features, we transform them into the same dimension using an affine transformation:
[…]
where $W$ and $b$ are learnable parameters. Then, we concatenate them to get the initial embedding representation of $x_{i}$:
(4) ${e}_{i}\in\mathbb{R}^{f\times d}=\left[{e}_{i1},{e}_{i2},\ldots,{e}_{if}\right],$
where $d$ is the dimension size of each field, and $f$ is the number of feature fields. The encoder processes input features using stacked multi-head self-attention and feed-forward layers:
(5) $H_{i}=(...FFN({SelfAttn}(e_{i}))).$
To standardize decoder training, we introduce the target sequence with a start-of-sequence token <SOS><SOS> and padding tokens <PAD><PAD>. These tokens do not represent any semantic value in the label space (i.e., $g(\{<SOS>,<PAD>\})=0$). The decoder autoregressively generates the target sequence based on the encoder’s output $H_{i}$ and the preceding subsequence. At training timestep $l$, the output token is
(6) $\hat{s}_{i}^{l}=\arg\max_{w\in\mathcal{V}}P_{\theta}\left(w\mid{H}_{i},\hat{s}_{i}^{<l}\right).$
Pre-training Optimization. During pre-training, we equip the model with fundamental market perception capabilities through supervised fine-tuning (SFT):
1. Cross-entropy loss for next-token prediction:
(7) ${L}_{NTP}=-\sum_{i=1}^{N}\sum_{l=1}^{L}\log P_{\theta}\left(\hat{s}_{i}^{l}\mid{H}_{i},\hat{s}_{i}^{<l}\right),$
where labels are derived from shading ratios generated by a two-stage method, we adopt a teacher forcing (TF) strategy to improve model efficiency, which directly feeds the ground truth $s_{i}^{l}$ as input at step $l+1$ to guide model training. We design a gate to enable the TF strategy adaptively:
(8) $TF_{pro}=TF_{pro}\cdot\frac{\eta}{\eta+e^{\left(\frac{iteration}{\eta}\right)}},$
where $TF_{pro}$ denotes the probability of employing the TF strategy, and $\eta$ is a hyperparameter. The gate mechanism decreases the probability of using TF as the number of training iterations increases, promoting faster convergence and improved model stability.
2. Negative Log-Loss of expected surplus for winning data $N_{+}$:
(9) $L_{\text{surplus}}=\sum_{i=1}^{N_{+}}\mathbb{E}((v_{i}-C(v_{i}*\alpha_{i}))\cdot Wr(v_{i}*\alpha_{i}\mid x_{i})),$
where $\alpha_{i}$ is the predicted shading ratio, $C(\cdot)$ is the cost function ($C(b)=b$ in FPA), and $Wr(\cdot)$ is the estimated winning rate (see Section 4.2.1). Notably, the argmax operation in Equation 6 obstructs gradient backpropagation. Therefore, we employ the gumbel softmax to obtain a differentiable distribution during token generation. At time step $l$, the dimension of the estimated raw output vector is $V$ (the size of the vocabulary). We do not consider the entire vocabulary space when calculating the weighted vector. Since the vocabulary is strictly ordered, we focus on the domain $V^{\prime}$, which corresponds to the neighborhood of the indices of the highest output values. For each position $p\in V^{\prime}$, we sample $u_{p}$ from the uniform distribution $U\sim\operatorname{Uniform}(0,1)$, and generate gumbel noise $n_{p}=-\log\left(-\log\left(u_{p}\right)\right)$ through the inverse transform, adding it to the original logits:
(10) $weight_{p}^{l}=\frac{\exp\left(\left(l_{p}+n_{p}\right)/\tau\right)}{\sum_{j=1}^{V^{\prime}}\exp\left(\left(l_{p}+n_{p}\right)/\tau\right)},$
where $l_{p}$ is the logit predicted by the decoder at step $l$ and position $p$, and $\tau$ is the temperature coefficient. As $\tau$ approaches 0, the output resembles discrete sampling; as $\tau$ approaches infinity, the output approximates a uniform distribution. Finally, we modify the decoding function $g(\cdot)$ using gumbel softmax weights and use the embedding of weighted sum $\hat{s}_{i}^{l}$ as input for the next token prediction. Compared to argmax, gumbel softmax not only addresses the issue of differentiability but also reduces the impact of errors in the earlier positions of the output sequence, increasing the stability.
In the pre-training stage, the loss is ${L}={L}_{NTP}+\lambda\cdot{L}_{surplus}$, where $\lambda$ is a hyperparameter.
4.2. Reward Preference Alignment System
Pre-trained models are limited by two-stage solutions and often get stuck in local optima due to the SFT in non-convex surplus loss and bias from training only on winning bids. To address this, we use reward-based preference alignment during post-training. With policy reinforcement learning, the model receives reward signals that help it develop a global perspective to escape local optima. Specifically, a surplus preference reward prevents overbidding, while an exploratory utility reward balances information and cost, reducing data selection bias. In this section, we first introduce the reward model and then introduce the two reward preference alignment systems.
Figure 4. The architecture of CHNet. It outputs the PDF distribution through the hierarchical dynamic network.
4.2.1. Channel-Aware Hierarchical Dynamic Network (CHNet)
Bid landscape forecasting is critical for evaluating bid quality. Following the previous works (Ren et al., 2019; Huang et al., 2024), we first discretize the continuous space via equal frequency binning, where a set of $T$ prices $0<b_{1}<b_{2}<\ldots<b_{T}$ covers the finite precision of price determinations. We have the probability density function (PDF) $P(z=b\mid x)$ of the winning price $z$ being $b$, and the cumulative distribution function (CDF) $Wr(b_{t}\mid x)=P(z\leq b_{t}\mid x)=\sum_{j\leq t}P(z\in B_{j})$ representing the winning rate of a certain bid $b_{t}$, where t=1,2,Tt=1,2,...T and $B_{j}=(b_{j},b_{(j+1)}]$ is a divided disjoint interval. Once the PDF curve of a request bid is obtained, the winning rate for another bid can be derived with a single additional CDF integral calculation.
In real-world advertising systems, varying ADXs/SSPs employ opaque and diverse billing methods, leading to significant differences in the bid landscape across channels. Inspired by multi-scenario approaches (Gao et al., 2023), we propose a channel-aware hierarchical dynamic network (CHNet) for fine-grained channel information modeling, as shown in Fig. 4. CHNet utilizes a hierarchical structure with an explicit channel-oriented layer and multiple implicit channel-oriented layers, sequentially modeling coarse-grained explicit information and fine-grained implicit information. For processing feature matrices222for simplification, request $i$ is not separately denoted in this section, we adhere to the method outlined in Section 4.1, with separate processing for channel feature $e^{c}$ and others $e^{o}$.
Explicit channel-oriented layer. We first perform global feature interactions for the common feature $e^{o}$ using a multi-layer perceptron (MLP), denoted as $e^{global}=MLP(e^{o})$. Furthermore, we adaptively generate the weight parameters based on the channel information $e^{c}$ via a re-parameterization approach, specifically by applying an MLP followed by a reshape operation: $W_{exp},b_{exp}=Reshape(MLP(e^{c}))$. These parameters are then used to instantiate the explicit scenario-oriented layer, resulting in the explicit channel representations:
(11) $e^{{explicit}}={FC}(e^{{global}};W_{exp},b_{exp}).$
Implicit channel-aware module. After explicit channel modeling, CHNet uncovers implicit patterns using a channel-aware multi-head attention mechanism. This enables comprehensive modeling of complex data distributions and adaptive identification of key implicit patterns. First, channel embeddings $e^{c}$ are processed by a MLP to generate raw weights, which are then normalized into $G$ groups ($G$ is the number of attention heads) using the softmax :
(12) $\begin{gathered}{weight}_{{ori}}={Reshape}\left({MLP}\left(e^{c}\right)\right),\\ {weight}_{{norm}}[g]={Softmax}\left(\text{weight}_{{ori}}[g]\right),g\in[1,G].\end{gathered}$
Next, we perform an element-wise product of the normalized weight with the feature embedding: $e^{weight}={weight}_{{norm}}\otimes{e}^{o}$.
Then, using the re-parameterization technique, we generate the weight parameters for different implicit channel layers, denoted as
(13) ${W}_{{imp}}[g],{b}_{{imp}}[g]={Reshape}({MLP}({e^{weight}}[g])),g\in[1,G].$
We input the explicit representations into each implicit channel-oriented layer for implicit modeling. After that, we concatenate the outputs of all implicit layers to predict the PDF distribution:
(14) $\begin{gathered}{e}_{g}^{{implicit}}={FC}(e^{{explicit}};W_{imp}[g],b_{imp}[g]),\\ output={Softmax}({FC}({Concat}({e}_{1}^{{implicit}},\dots,{e}_{G}^{{implicit}}))).\end{gathered}$
Training loss. During the CHNet training process, the loss function is comprised of three components. For winning bid logs in a non-FPA scenario, we directly compute the negative log-likelihood loss based on the actual winning price: ${L}_{win}=-\sum_{i=1}^{N_{+}}\log p\left(c_{i}\mid x_{i}\right)$. Note that in the FPA scenario, this loss cannot be applied because the cost paid does not reflect the minimum winning price. For losing bid logs, we only know that the winning price exceeds the current bid, and we maximize the CDF of the right half of the bid, resulting in ${L}_{lose}=-\sum_{i=1}^{N_{-}}\log[1-Wr(b_{i}\mid x_{i})]$. Additionally, to further enhance the model’s ranking and classification capabilities (Lin et al., 2024), we introduce a pairwise loss related to the winning rate, denoted as
(15) ${L}_{{ranknet}}=-\frac{1}{N_{+}N_{-}}\sum_{i=1}^{N_{+}}\sum_{j=1}^{N_{-}}\log(\sigma(Wr(b_{i}|x_{i})^{(+)}-Wr(b_{j}|x_{j})^{(-)})).$
Finally, the loss function of CHNet is formulated as
(16) ${L}=\lambda_{1}\cdot{L}_{win}+\lambda_{2}\cdot{L}_{lose}+\lambda_{3}\cdot{L}_{ranknet}.$
Figure 5. The architecture of post-training with GRPO.
4.2.2. Surplus Optimized Alignment
The primary objective of bid shading is to prevent incurring excessive costs. Accordingly, we use the expected surplus associated with the shaded bid b=valphab=v*alpha as the reward $r_{surplus}=(v-C(b))\cdot Wr(b\mid x)$ and optimize the model using reinforcement learning strategies. We employ GRPO to optimize the strategy based on relative rewards within the group. More specifically, for each bid request reqreq, GRPO samples a group of outputs $\{\alpha_{1},\ldots,\alpha_{2},\ldots,\alpha_{m}\}$ from the old policy and then optimizes the policy model by maximizing the following objective:
(17) $\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{req\sim P(R),\{\alpha_{i}\}_{i=1}^{m}\sim\pi_{\theta_{\text{old}}}}\Bigg{[}\frac{1}{m}\sum_{i=1}^{m}\min\bigg{(}\frac{\pi_{\theta}(\alpha_{i}\mid req)}{\pi_{\theta_{\text{old}}}^{\prime}(\alpha_{i}\mid req)}A_{i},$
$\operatorname{clip}\Big{(}\frac{\pi_{\theta}(\alpha_{i}\mid req)}{\pi_{\theta_{\text{old}}}^{\prime}(\alpha_{i}\mid req)},1-\epsilon,1+\epsilon\Big{)}A_{i}\bigg{)}-\beta\cdot\mathbb{D}_{\text{KL}}\big{[}\pi_{\theta}\|\pi_{\text{ref}}\big{]}\Bigg{]},$
where $A_{i}$ denotes the advantage, which is calculated solely based on the relative return of each group’s internal output:
(18) $A_{i}=\frac{r_{i}-\operatorname{mean}(\{r_{1},r_{2},...,r_{m}\})}{\operatorname{std}(\{r_{1},r_{2},...,r_{m}\})}.$
Additionally, we add the KL divergence between the training and reference strategies in the objective. GRPO simultaneously employs the group-relative advantage estimation and a KL divergence penalty to ensure that strategy updates are efficient and stable. The complete algorithm processing is detailed in Algorithm 2.
Pre-training with surplus-based SFT restricts the model to exploring the specific output bid and may cause gradient descent to get stuck in local optima due to the non-convex surplus. In contrast, policy reinforcement learning expands the generation space for getting global optima. Additionally, since the bidding process is uncertain and each winning price reflects a probability distribution, we incorporate losing bid data with a smaller weight during post-training.
4.2.3. Exploration Utility Alignment
In the DSP advertising bidding scenario, data selection bias presents a significant challenge, as the collected data are generated using a fixed bidding algorithm, which complicates the bid landscape forecasting. To address this issue, we incorporate bid exploration into the bid shading system, providing more robust data support for maximizing long-term surplus.
According to uncertainty sampling in active learning (Settles and Craven, 2008; Wang et al., 2024), we think that the greater the predicted uncertainty, the more informative the exploration will be. We quantify the exploration utility using the entropy of the winning rate. For a bid request $x_{i}$, the predicted winning rate $Wr_{i}=Wr(b_{i}\mid x_{i})$ at a particular bid $b_{i}$ can be framed as a binary classification problem (win or lose). Thus, the entropy of the bid request at bid $b_{i}$ is denoted as
(19) $\phi^{E}(x_{i},b_{i})=-(Wr_{i}\cdot log(Wr_{i})+(1-Wr_{i})\cdot log(1-Wr_{i})).$
Additionally, during exploration, it is important to avoid excessive cost increases; the final exploration reward is denoted as $r_{exploration}=\frac{\phi^{E}(x_{i},b_{i})}{C(b_{i})}$. When calculating the final advantage $A_{i}$ (Equation 18), the reward is
(20) $r=r_{surplus}+\omega\cdot r_{exploration},$
where $\omega$ is not a fixed parameter. It is unnecessary to explore all samples; by identifying samples with high similarity, we can avoid unproductive exploration of outliers. Inspired by the metric learning (Snell et al., 2017), we compute the exploration weight for request $x_{i}$:
(21) $\omega_{i}=cos(e_{i},\frac{1}{\left|N\right|}\sum_{j=1}^{N}e_{j}),$
where cos(·, ·) is the cosine similarity function, more representative requests are assigned greater exploration weight. The post-training structure is shown in Figure 5.
5. Experiments
In this section, we conduct extensive experiments to verify the performance of GBS, where implementation details are detailed in Appendix C.
5.1. Experimental Setup
5.1.1. Datasets
We evaluate our method on both the public and real-world industrial datasets, with detailed statistics presented in Table 1. The public iPinYou (Liao et al., 2014) data are derived from the SPA setting, where the advertiser’s bid is used as the actual value in the bid request, and the paid price is regarded as the winning price. The dataset is randomly split into training and testing sets with a ratio of 7/3. Our private dataset is collected from one month of bid sampling data from Meituan DSP and is partitioned similarly. Notably, only the testing set contains minimum winning price, which is used for evaluation.
Table 1. The statistics of datasets.
DataSet Size Winning rate features
iPinYou 10,577,061 29.7% 18
Private 162,549,577 3.89% 197
5.1.2. Evaluation Metrics
For offline evaluation in FPA, business performance is crucial in determining whether an algorithm is suitable for production deployment, as it reflects the algorithm’s effectiveness. Given the objective of bid shading, we use surplus $(V-b)\cdot I(b>z)$ and surplus rate as key business metrics, as they directly indicate the impact of the algorithm on business outcomes. Specifically, we calculate the surplus rate as the percentage of residual surplus relative to the total optimal surplus:
(22) $\text{surplus rate (SR)}=\frac{\sum_{i}(v_{i}-b_{i})I(b_{i}>{z}_{i})}{\sum_{i}(v_{i}-z_{i})}.$
5.2. Overall Performance
We compare the performance of GBS with several state-of-the-art baselines, including the bid shading and the generative methods:
EDDN (Zhou et al., 2021): A two-stage bid shading method, which optimizes the surplus using golden section search in the OR stage.
WR (Pan et al., 2020): A two-stage bid shading method, which optimizes the surplus using bisection algorithm in the OR stage.
TSBS-DLF (Ren et al., 2019): A two-stage bid shading method uses the DLF model in the ML stage, which models the bid landscape without distributional assumptions, employing the conditional probability chain rule and LSTM.
TSBS-ADM (Li et al., 2022): A two-stage bid shading method uses the ADM model in the ML stage, which uses neighborhood likelihood loss for accurate prediction.
MEBS (Gong et al., 2023): An end-to-end bid shading method, which jointly optimizes the shading model and the win rate model, performs supervised learning through the negative logarithm of the surplus as the loss.
CVAE (Sohn et al., 2015): An extended model of variational autoencoder, which achieves sample generation based on specific inputs by integrating conditional variables in the encoder and decoder. We optimize it through the pre-training method proposed in this paper.
DF (Ho et al., 2020): A generative diffusion model that corrupts the data distribution by gradually adding noise and then learns the inverse denoising process to generate shading ratio. We optimize through the pre-training method proposed in this paper.
Post-CVAE & Post-DF: Post-training will be performed on the pre-trained generative model, which uses reparameterization to transfer gradients.
Based on the results in Table 2, we can draw the following conclusions: 1) The traditional two-stage algorithm performs significantly worse on the Meituan private dataset, which suggests that a more complex market environment poses greater challenges for bid shading, and that coarse-grained bid landscape forecasting and binary search cannot adapt to a broader range of business scenarios; 2) The traditional generative model shows varying degrees of improvement after post-training and outperforms the two-stage method, indicating that policy reinforcement learning can effectively enhance these models to raise their performance ceiling, and the end-to-end generative approach offers distinct advantages; 3) GBS achieves outstanding results across all metrics. By combining an autoregressive residual generative model with post-training, GBS effectively mitigates the suboptimality of traditional methods and adapts well to complex market environments.
Table 2. Overall offline experiment results.
Model iPinYou Private
SR Surplus SR Surplus
EDDN 47.82% 90,199,768 19.88% 14,582,074
WR 54.90% 103,560,851 26.13% 19,164,882
TSBS-DLF 55.22% 104,164,484 27.25% 19,986,132
TSBS-ADM 55.49% 104,673,800 32.67% 24,183,220
MEBS 54.46% 102,716,112 31.45% 23,041,120
CAVE 51.42% 96,981,112 36.45% 26,698,064
Post-CAVE 57.14% 107,781,128 36.63% 26,830,032
DF 48.04% 90,617,280 31.28% 22,912,038
Post-DF 51.13% 96,445,909 33.89% 24,823,816
GBS 60.48% 114,068,392 41.74% 30,580,080
5.3. Ablation Study
To assess the effectiveness of each component in GBS, we conducted a series of ablation studies. Specifically, we build several variants of the GBS:
w/o PostT: A variant of GBS without the Post-training stage, which directly uses the pre-trained model to output shading ratio.
w/o PostT-EUA: A variant of GBS without the exploration utility alignment in post-training, which does not perform bid exploration and only performs surplus optimization.
w/o GM: A variant of GBS without the autoregressive generative model, which uses the traditional generative model CVAE (Sohn et al., 2015) to generate the shading ratio and keeps the pre-training and post-training consistent.
w/o CHNet: A variant of GBS without the CHNet, which using a binary classification model to provide the predicted winning rate.
Table 3 presents the results of the ablation study. CVAE is less effective than our GM, underscoring the importance of progressively outputting results through residuals in refined bid generation. Excluding CHNet and bid exploration also leads to similar performance drops, indicating that a more fine-grained reward model and a broader range of training data are crucial for enhancing model performance. Omitting post-training results in the most significant performance degradation, highlighting the limitations of relying solely on supervised learning for the bid shading task. The results suggest that policy reinforcement learning plays a vital role in helping the model overcome performance bottlenecks.
Table 3. The contributions of different components of GBS.
Model iPinYou Private
SR Surplus SR Surplus
w/o PostT 56.22% 106,036,888 36.91% 27,041,464
w/o PostT-EUA 60.16% 113,458,768 41.54% 30,429,570
w/o GM 57.14% 107,781,128 36.63% 26,830,032
w/o CHNet 58.22% 109,805,915 38.16% 27,953,616
GBS 60.48% 114,068,392 41.74% 30,580,080
5.4. Performance Analysis of CHNet
In this section, we evaluate the performance of the proposed CHNet in bid landscape forecasting on private datasets. We use the area under the curve (AUC) and binary cross-entropy (BCE) of winning rate as evaluation metrics. In the experiments, we compare the following competitive bid landscape forecasting methods: 1). Censorship Linear Model (CLM) uses the normal distribution as the prior for single-point estimation of the winning rate; 2). Deep Landscape Forecasting (DLF) model (Ren et al., 2019); 3). Arbitrary Distribution Model (ADM) (Li et al., 2022); 4). CHNet-w/o EUA is the ablation experiment without the exploration utility alignment module during the post-training phase.
Figure 6. The performance of CHNet.
The experimental results are shown in Fig. 6. The results demonstrate that CHNet performs superior ranking and classification tasks, owing to its refined modeling of channel information, a key characteristic of the market environment. Additionally, removing the bid exploration module decreases performance and exposes the model to data selection bias. Therefore, incorporating exploration in the bidding process is essential for improving the model’s generalization ability.
Figure 7. Architecture of the online deployment with GBS.
5.5. Performance on Online System (Meituan)
To evaluate the online performance of GBS, we deployed it on the Meituan DSP. We conducted a rigorous A/B test over two weeks, and the engineering architecture is shown in Fig. 7. We allocated 30% of the traffic to GBS, while the baseline method was an online two-stage bid shading approach, which utilizes the golden section search in the operations research stage. We model certain channels as a non-ideal SPA scenario, consistent between baseline and GBS. We use four key online metrics for evaluation: Return On Investment (ROI), Cost Per Mille (CPM), Cost Per Click (CPC), and Inference Time (IT).
Table 4 shows the improvement of GBS relative to the baseline. The results indicate that GBS has significantly improved in all online metrics compared to the baseline, achieving significant business benefits. During the inference stage, GBS directly outputs the shading ratio, resulting in lower inference time than the two-stage method and reducing the burden on online services.
Table 4. Online A/B test results.
ROI CPM CPC IT
Baseline (Two-stage) 0.0% 0.0% 0.0% 0.0%
GBS +3.4% -4.1% -4.9% -37.6%
6. Conclusion and Future Work
This paper proposes a novel Generative Bid Shading (GBS) framework that addresses the suboptimality of existing bid shading techniques. GBS uses an autoregressive generative model to decompose the shading ratio into conditionally dependent sequences and incorporates a channel-aware reward model to capture fine-grained bid landscape features. By integrating the reward preference alignment system with the GRPO, GBS prevents overbidding, mitigates sample selection bias, and enables efficient policy updates. Extensive experiments show that GBS outperforms state-of-the-art models.
In the future, we will further optimize the generative model to better adapt upstream ranking scores and explore additional reward systems with advanced optimization algorithms.
Appendix A Vocabulary Construction
We use the shading ratios generated by the two-stage method as the base dataset to construct the vocabulary. The process (Algorithm 1) is initialized with a high starting quantile $q_{start}$, and then adaptively reduced by a decay rate $\delta$ until reaching the terminal quantile $q_{end}$.
1
Require: Shading ratios from two stage method $Y=\{\alpha_{j}\}_{j=1}^{N}$, initially empty Vocabulary $V=\{\}$, start percentile $q_{{start}}$, end percentile $q_{{end}}$, decay rate $\delta$ , minimal restoration error $\epsilon_{1}$ and $\epsilon_{2}$.
2
3Sort $Y$ in descending order to obtain $\hat{Y}=\{\hat{\alpha}_{j}\}_{j=1}^{N}$;
4 Initialize iteration counter $i=1$, error metric $err=\infty$, current percentile $q=q_{{start}}$;
5
6while $err>\epsilon_{1}$ do
7   Compute the $q$-percentile $o_{i}$ of $\hat{Y}$;
8
9 if $o_{i}\leq\epsilon_{2}$ then
10    break;
11
12   end if
13
14  Generate a new token $w_{i}$ which satisfy $o_{i}=g(w_{i})$ and insert $w_{i}$ into vocabulary $V$;
15
16  Update $\hat{Y}$ using:
$\hat{y}_{j}=\begin{cases}\hat{y}_{j},&\text{if}\ \hat{y}_{j}<o_{i},\\ \hat{y}_{j}-o_{i},&\text{otherwise}\end{cases}$
17  Update the error metric:
$err=\max\left\{\frac{\hat{y}_{j}}{y_{j}}\right\}_{j=1}^{N}$
18  Update percentile $q$ with decay rate $\delta$:
$q=\max(q\cdot\delta,q_{{end}})$
19  Increase $i$: $i\leftarrow i+1$;
20
21 end while
22
Result: Vocabulary $V$
Algorithm 1 Constructing Vocabulary with Dynamic Percentiles
Appendix B Post-training process of GBS
The post-training process is detailed in Algorithm 2.
1
Require: Pre-trained generative model $\pi_{\theta_{\text{init}}}$, reward system, post-training data $\mathcal{D}$
2
3Initialize policy model $\pi_{\theta}\leftarrow\pi_{\theta_{\text{init}}}$;
4 for iteration =1=1 to II do
5   Update reference model $\pi_{\text{ref}}\leftarrow\pi_{\theta}$;
6 for step =1=1 to $S$ do
7      Sample batch $\mathcal{D}_{b}\leftarrow\mathcal{D}$;
8      Set old policy $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$;
9    foreach bid request $req\in\mathcal{D}_{b}$ do
10         Sample $m$ outputs $\{\alpha_{i}\}_{i=1}^{m}\sim\pi_{\theta_{\text{old}}}(\cdot\mid req)$;
11         Compute final rewards $\{r_{i}\}_{i=1}^{m}$ by Equation 20;
12         Estimate token advantages $\hat{A}_{i,t}$ via group-relative method (Equation 18);
13         Update $\pi_{\theta}$ by maximizing GRPO objective (Equation 17);
14
15      end foreach
16
17   end for
18
19 end for
20
Result: optimized policy model $\pi_{\theta}$
Algorithm 2 Post-training Stage: GRPO Optimization
Appendix C Implementation Details
We implement all deep learning baselines and the proposed GBS on NVIDIA A100-SXM4-80GB GPUs. During the pre-training stage, we use the Adam optimizer with a batch size of $2^{10}$ and a learning rate of $1e^{-4}$. The maximum output sequence length $L$ is set to 5, $\eta$ in the TF Gate is 5000, and the temperature coefficient $\tau$ in gumbel softmax is 0.2. In the post-training stage, the loss hyperparameters $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ for CHNet training are set to 0.5, 1, and 0.2, respectively, while the KL weight $\beta$ in the GRPO optimization function is set to 0.01. We divide the training and testing sets of the iPinYou and private datasets into 100 segments each to simulate the continuous arrival of bid requests in an online environment. We calculate the surplus by the winning price in each round and collect the win/loss labels. Subsequently, the bid data are used to update the CHNet and the generative model. This process is repeated for 100 rounds until the experiment is complete.